# Impact of Fixing Spins in a Quantum Annealer with Energy Rescaling

### v1

@@ Abstract
Quantum annealing is a promising algorithm for solving combinatorial optimization problems. However, various hardware restrictions significantly impede its efficient performance. Size-reduction methods provide an effective approach for addressing large-scale problems but often introduce additional challenges. A notable hardware restriction is the limited number of decision variables quantum annealing can handle compared to the size of the problem. Moreover, when employing size-reduction methods, the interactions and local magnetic fields in the Ising model––used to represent the combinatorial optimization problem––can become excessively large, making them difficult to implement on hardware. Although prior studies suggest that energy rescaling impacts the performance of quantum annealing, its interplay with size-reduction methods remains unexplored. This study examines the relationship between fixing spins, a promising size-reduction method, and the effects of energy rescaling. Numerical simulations and experiments conducted on a quantum annealer demonstrate that the fixing spins method enhances quantum annealing performance while preserving the spin-chain embedding for a homogeneous, fully connected ferromagnetic Ising model.
@@ 1 Introduction
In addition to size restrictions, the hardware limitations associated with graph structures in quantum annealers contribute to suboptimal solutions. Ising models with denser graphs than those supported by the quantum annealer’s architecture cannot be faithfully represented. To bridge this gap, embedding methods have been developed [28, 29]. These techniques utilize qubit copies, referred to as spin chains, to effectively represent the couplers of dense Ising models. While such embeddings allow quantum annealers to handle dense graphs, they require a significant number of qubits and demand careful parameter tuning, particularly the chain strength between qubit copies. Setting optimal parameters is a challenging task, making embedding one of the factors that can compromise solution quality.
The second approach seeks to minimize the embedding overhead by reducing the problem size. In dense graph scenarios, the number of required couplings grows significantly with the number of spins. For instance, a fully connected graph with $n$ spins requires $n-1$ couplings, demanding graph structures with higher connectivity as the number of spins increases.
The third approach involves error mitigation techniques. Previous studies have explored methods to relax errors and maintain solution accuracy by parallelizing the graph structure [30, 31, 32, 33]. Collectively, these approaches demonstrate diverse strategies to address the graph structure constraints of quantum annealers. Rescaling techniques are employed to embed the Ising model onto a practical quantum annealer, overcoming limitations related to energy scales. Recent studies suggest that energy rescaling impacts noise strength [34]. Leveraging this property, researchers have estimated ideal energy parameters using zero-noise extrapolation. Another study treated rescaling parameters as variational parameters, showing that optimized rescaling significantly enhances performance [35]. Therefore, incorporating rescaling considerations is essential for achieving high-quality solutions.
@@ 2 Setup
Here, $s\in[0,1]$ denotes the normalized time, where the annealing time $\tau$ relates to the normalized time as $s=\frac{t}{\tau}$, with $t$ being the elapsed time.
@@
The selection of spins to be fixed, along with their respective directions, is determined using a classical pre-processing method from previous studies [22, 23, 24, 25, 26]. In this study, to characterize the fixed spins without assuming a predetermined selection, we adopt the error probability $p_{\mathrm{err}}$ as defined in the previous study [21]. The error probability $p_{\mathrm{err}}$ is the proportion of spins that are aligned with the direction of the ground state of the Ising model within the spin configuration. The previous study shows the effectiveness of fixing spins for large-scale combinatorial optimization problems in quantum annealer. Also, fixing spin method enhance the performance of the quantum annealing from the perspective of energy gap. However, the impact of the error of fixing spins on the energy gap are not revealed.
@@
The rescaling parameters are affected by fixing spins because the interaction terms $J_{ij}\sigma_{i}^{z}\sigma_{j}^{z}$ are transformed into $J_{ik}\sigma_{i}^{z}s_{k}$, which function as effective local magnetic fields, as shown in Eq. (6). Consequently, the fixing spins method alters the range of local magnetic fields, thereby influencing the rescaling parameters. This study examines the relationship between the fixing spins method and rescaling. A previous study [21] demonstrated performance improvements through fixing spins. However, this conclusion may not hold when rescaling is considered.
@@ 3 Results
Figure 2: (Color online) The color map shows the distribution of the minimum energy in the $n$ versus $r$ space for the homogeneous fully connected ferromagnetic Ising model with $N=160$. $n=160$ corresponds to the results without fixed spins. The annealing time is set to $\tau=2000.0~{}\mu s$. The ground state energy is $0.597$. Each panel represents a different error rate for fixing spins, with $p_{\mathrm{err}}$ set to (a) $p_{\mathrm{err}}=0.0$, (b) $p_{\mathrm{err}}=0.1$, (c) $p_{\mathrm{err}}=0.2$, (d) $p_{\mathrm{err}}=0.3$, (e) $p_{\mathrm{err}}=0.4$ and (f) $p_{\mathrm{err}}=0.5$, respectively.
@@
The normalization by $1/N$ ensures that $\mathcal{H}_{\mathrm{p}}$ is an extensive quantity. If the ground state is degenerate, determining the error rate becomes challenging because the Hamming distance from the ground state is not uniquely defined. Therefore, we consider the Ising model without degeneracy. The ferromagnetic Ising model with $h=0$ exhibits trivial twofold degeneracy. To eliminate this trivial degeneracy, we set $h=0.1$ and $J=1$. The ground state of this homogeneous fully connected Ising model is trivially the all-up state.
The D-Wave Advantage system 6.4 [37] was used as the quantum annealer for the experiment. This system features $5760$ physical qubits arranged in the Pegasus graph [38], enabling the embedding of a complete graph with $177$ logical qubits through minor embedding [28, 29]. However, not all qubits in the system are functional. Consequently, the number of spins in the Ising model was set to $N=160$, the near maximum size that can be embedded in the quantum annealer.
The Ising model is embedded in the quantum annealer using minor embedding, and the chain strength—representing the interaction between physical qubits—must be adjusted to ensure that spins within the same spin chain align in the same direction. If the auto-scale method is employed [39], the chain strength increases with the length of the spin chain. As a result, the chain strength must remain within the range $[\mathcal{J}_{\mathrm{min}},\mathcal{J}_{\mathrm{max}}]$ for the interaction. For this study, the chain strength was set to the maximum value of $2.0$, which is the strongest value permissible in this configuration[39].
The ground state search in the quantum annealer was performed $100$ times due to the stochastic behavior of the system, and the minimum energy from these $100$ runs was adopted as the energy output of the quantum annealer in this experiment.
Additionally, we investigated the energy gap using exact diagonalization for the same fully connected Ising model, without considering the embedding effect. The energy gap analysis yielded similar results for the long annealing time limit. Finally, we investigated the properties of the system at the thermodynamic limit, accounting for the effects of rescaling and fixing spins.
@@ 3.1 Experiment on an actual quantum annealer
When embedding an Ising model with a large energy scale into a real quantum annealer, rescaling the energy scale becomes unavoidable. This rescaling is performed as described in Eq. (2). The experiment was conducted on the D-Wave Advantage system 6.4 to examine the dependence of the quantum annealer’s performance on the rescaling parameter.
Figure 2 illustrates the parameter dependence of the energy. Since the energy scale after rescaling differs from the pre-rescaling energy scale, the energy values are restored to the original Ising model, as expressed in Eq. (11), after solving them with the quantum annealer. The $x$-axis represents the number $n$ of spins after fixing spins, and the $y$-axis denotes the rescaling parameter $r$, as shown in Eq. (9). As $r$ decreases, the quantum annealer achieves lower energy, as shown in Fig 2 (a)-(c). This result indicates that a wide parameter range is necessary to maintain the performance of the quantum annealer.
@@
Eq. (12) exhibits the maximum increase in the range of the local magnetic field due to fixed spins. Since we are dealing with a homogeneous fully connected ferromagnetic Ising model, after fixing spins with $p_{\mathrm{err}}=0$, the local magnetic fields correspond to this upper bound, as shown in Eq. (12). When fixing spins with $p_{\mathrm{err}}=0$, the local magnetic fields change as follows:
$h^{\prime}=h+J(N-n)=h+\sum_{k=n+1}^{N}|J_{k}|.$ (13)
@@
High energy points observed in Fig. 2 (d)-(f) result from the breaking of the embedding chain in the hardware. The local magnetic field is distributed equally among the spins within the spin chain. Consequently, the local magnetic fields are weaker compared to before the spin chain was constructed. Consequently, a longer chain length requires a stronger chain strength to prevent chain breakage. As the number of spins increases, a stronger spin chain is needed to maintain its integrity. For a spin chain to remain intact when the chain strength is weak, the spins in the chain must have identical values, without the interaction effects between them.
Therefore, with a small $h$, it becomes increasingly difficult to maintain the spin chain as its length grows. Consequently, with large $n$ and $p_{\mathrm{err}}$, the spin chain is more likely to break compared to other parameter ranges. Additionally, because the Ising model parameters are extensive and the chain strength is relatively weak, the chain breaks more frequently when $r$ is small. To prevent chain breakage, expanding the parameter range is necessary to accommodate large-scale problems when embedding into the quantum annealer. These results highlight the importance of fixing spins in the correct direction to achieve lower energy and preserve spin chains on the hardware.
@@
Figure 4: (Color online) Color map of the minimum energy gap $\Delta_{\mathrm{min}}$ distribution in $n$ versus $r$ space for the homogeneous fully connected ferromagnetic Ising model with $N=160$. $n=160$ represents the results without fixing spins. Each panel corresponds to a different error rate $p_{\mathrm{err}}$ for fixing spins: (a) $p_{\mathrm{err}}=0.0$, (b) $p_{\mathrm{err}}=0.1$, (c) $p_{\mathrm{err}}=0.2$, (d) $p_{\mathrm{err}}=0.3$, (e) $p_{\mathrm{err}}=0.4$, (f) $p_{\mathrm{err}}=0.5$.
@@ 3.2 Energy gap analysis
Since the minimum energy gap between the ground state and the first excited state, $\Delta_{\mathrm{min}}$, reflects the performance of quantum annealing in the adiabatic limit, as per the adiabatic condition [36, 40], we investigated its properties. Figure 3 shows $\Delta_{\mathrm{min}}$ of the Hamiltonian after fixing spins with $p_{\mathrm{err}}=0$ and rescaling with $r$.
The number of fixing spins $n$ at which $\Delta_{\mathrm{min}}$ is maximized varies for each $r$ as shown in Fig. 3. Energy gap $\Delta_{\mathrm{min}}$ increases with $n$, consistent with a previous study [21]. However, the fixing spins method does not necessarily lead to an expansion of $\Delta_{\mathrm{min}}$ when rescaling is considered. For example, when $r=5/2$, the maximum $\Delta_{\mathrm{min}}$ occurs at $n=140$. These results indicate a trade-off between the energy gap expansion due to fixing spins and the reduction of the energy gap caused by rescaling. Additionally, for $r=1/4$, $\Delta_{\mathrm{min}}$ decreases as $n$ increases. This demonstrates that the effects of fixing spins depend on the rescaling parameter. The findings emphasize the importance of exploring the relationship between fixing spins and rescaling.
Next, we investigated the behavior of $\Delta_{\mathrm{min}}$, including errors in fixing spins. Figure 4 presents a heatmap of $\Delta_{\mathrm{min}}$ for different values of $r$ and $n$. Each panel corresponds to a different $p_{\mathrm{err}}$. As shown in Figure 4, as $p_{\mathrm{err}}$ increases or $r$ decreases, the optimal $n$ that maximizes $\Delta_{\mathrm{min}}$ becomes smaller. These results suggest that both $r$ and $n$ must be carefully considered to achieve high-quality solutions. For $r=1.0$, fixing spins (lower $n$) improves $\Delta_{\mathrm{min}}$. In contrast, for $r=10$, fixing spins reduces $\Delta_{\mathrm{min}}$. These findings highlight the necessity of considering the rescaling parameters when fixing spins and adjusting parameters accordingly. Fixing spins has a positive effect at small $r$, but a negative effect at large $r$.
@@
$\Delta(s)$ =$=\left\{s^{2}\sin^{2}{\theta_{0}}\left(\frac{h}{r}+\frac{1+\sin^{2}{\theta_{0}}}{s\sin{\theta}}\right)^{2}-\frac{(1+\sin^{2}{\theta_{0}})^{2}}{s^{2}\sin^{2}{\theta}}\right.$
$+\left.s^{2}\sin^{4}{\theta_{0}}\frac{J^{2}}{r^{2}}+s[2\sin^{2}{\theta_{0}}-(1-s)\cos^{3}{\theta_{0}}]\frac{J}{r}\right.$
$\left.+(1-s)^{2}\cos^{2}{\theta_{0}}\right\}^{\frac{1}{2}},\quad 0\leq s\leq 1.$ (15)
@@
In Fig. 5, $\Delta_{\mathrm{min}}$ increases as $r$ decreases, in line with the increase in the local magnetic fields $h^{\prime}$ after fixing spins. This indicates that smaller values of $r$ are more favorable for quantum annealing, assuming no restrictions on the parameter range. The descending order of $h^{\prime}$ corresponds directly to the descending order of $\Delta_{\mathrm{min}}$. However, it is important to note that the allowed parameter range differs for each $r$ in Fig. 5, even though the parameter range is constrained in actual hardware. Therefore, the comparison is more advantageous for settings with smaller $r$.
We also investigated $\Delta_{\mathrm{min}}$ while considering the restrictions of the parameter range. The black lines in Fig. 5 illustrate a comparison of $\Delta_{\mathrm{min}}$ within the same $h$ range $[\mathcal{M}_{\mathrm{min}},\mathcal{M}_{\mathrm{max}}]$, where we assume $|\mathcal{M}_{\mathrm{min}}|=|\mathcal{M}_{\mathrm{max}}|$. We focus solely on the $h$ range because, in this numerical simulation, the $h$ range is always broader than the $J$ range after fixing spins. According to Fig. 5, the point at which the maximum energy gap occurs in the line of $|\mathcal{M}_{\mathrm{min}}|=0.08$ is at $r=1/2$, while for the line of $|\mathcal{M}_{\mathrm{min}}|=0.8$, it occurs at $r=1.0$. These results suggest that optimal rescaling parameters exist for different $h$ ranges. Therefore, the effects of fixing spins in the thermodynamic limit also depend on the rescaling parameter.
When the $h$ range is wide, the fixing spins method is effective in expanding $\Delta_{\mathrm{min}}$. In contrast, when the $h$ range is narrow, the fixing spins method becomes less effective because the rescaling effects, which reduce $\Delta_{\mathrm{min}}$, dominate over the benefits of expanding $\Delta_{\mathrm{min}}$ through fixing spins. These findings are also supported by the finite-size Ising model results in Fig. 4. The same considerations apply to the energy gap at finite sizes, even if the problem scale increases. Therefore, the properties of the energy gap with fixing spins and rescaling are extensive with respect to size.

### v3

@@ Abstract
Quantum annealing is a promising algorithm for solving combinatorial optimization problems. However, various hardware restrictions significantly impede its efficient performance. Size-reduction methods provide an effective approach to addressing large-scale problems but often introduce additional difficulties. A notable hardware restriction is that quantum annealing can handle only a limited number of decision variables, compared to the size of the problem. Moreover, when employing size-reduction methods, the interactions and local magnetic fields in the Ising model––used to represent the combinatorial optimization problem––can become excessively large, making them difficult to implement on hardware. Although prior studies suggest that energy rescaling impacts the performance of quantum annealing, its interplay with size-reduction methods remains unexplored. This study examines the relationship between fixing spins, a promising size-reduction method, and the effects of energy rescaling. Numerical simulations and experiments conducted on a quantum annealer demonstrate that the fixing spins method enhances quantum annealing performance while preserving the spin-chain embedding for a homogeneous, fully connected ferromagnetic Ising model.
@@ 1 Introduction
In addition to size restrictions, the hardware limitations associated with graph structures in quantum annealers contribute to suboptimal solutions. Ising models with denser graphs than those supported by the quantum annealer’s architecture cannot be faithfully represented. To bridge this gap, embedding methods have been developed [28, 29]. These techniques utilize qubit copies, referred to as qubit chains, to effectively represent the couplers of dense Ising models. While such embeddings allow quantum annealers to handle dense graphs, they require a significant number of qubits. Here, the length of the qubit copies is called chain length. The qubit chains and demand careful parameter tuning, particularly the chain strength between qubit copies. Setting optimal parameters is a challenging task, making embedding one of the factors that can compromise solution quality.
Several approaches have been proposed to alleviate the graph structure restrictions in quantum annealers. The first approach focuses on reducing the number of spins required for embedding by increasing the graph density on the hardware. This effort has led to the development of increasingly dense graph architectures [30].
The second approach seeks to minimize the embedding overhead by reducing the connectivity between qubits by reducing the problem size. In dense graph scenarios, the number of required couplings grows significantly with the number of qubits. For instance, a fully connected graph with $n$ spins requires $n-1$ couplings, demanding graph structures with higher connectivity as the number of qubits increases.
The third approach involves error mitigation techniques. Previous studies have explored methods to mitigate errors and maintain solution accuracy by parallelizing the graph structure [31, 32, 33, 34]. Collectively, these approaches demonstrate diverse strategies to address the graph structure constraints of quantum annealers. Rescaling techniques are also employed to embed the Ising model onto a practical quantum annealer, overcoming limitations related to energy scales. Recent studies suggest that energy rescaling impacts noise strength [35]. Leveraging this property, ideal energy parameters have been estimated using zero-noise extrapolation. Another study treated rescaling parameters as variational parameters, showing that optimized rescaling significantly enhances performance [36]. Therefore, incorporating rescaling considerations is essential for achieving high-quality solutions.
@@ 2 Setup
Here, $s\in[0,1]$ denotes the normalized time, where the annealing time $\tau$ relates to the normalized time as $s=\frac{t}{\tau}$, with $t$ being the elapsed time. The annealing schedules $A(s)$ and $B(s)$ are monotonically decreasing and increasing functions, respectively, satisfying $A(0)\gg B(0)$ and $A(1)\ll B(1)$.
@@
The selection of spins to be fixed, along with their respective directions, is determined using a classical pre-processing method from previous studies [22, 23, 24, 25, 26]. In this study, to characterize the fixed spins without assuming a predetermined selection, we adopt the error probability $p_{\mathrm{err}}$ as defined in the previous study [21]. The error probability $p_{\mathrm{err}}$ is the proportion of spins that are aligned with the opposite direction of the ground state of the Ising model within the spin configuration. The previous study shows the effectiveness of fixing spins for large-scale combinatorial optimization problems in the quantum annealer. Also, the fixing spin method enhances the performance of the quantum annealing from the perspective of the energy gap. However, the impact of the error of fixing spins on the energy gap is not revealed. In this study, the error probability is easy to calculate since a homogeneous fully connected ferromagnetic Ising model, which has the trivial ground state, is used.
@@
The rescaling parameters are affected by fixing spins because the local magnetic fields are changed as shown in Eq. (6). Consequently, the fixing spins method alters the range of local magnetic fields, thereby influencing the rescaling parameters. This study examines the relationship between the fixing spins method and rescaling. A previous study [21] demonstrated performance improvements through fixing spins. However, this conclusion may not hold when rescaling is considered.
@@ 3 Results
Figure 2: (Color online) The color map shows the distribution of the minimum energy in the $n$ versus $r$ space for the homogeneous fully connected ferromagnetic Ising model with $N=160$. The spacing of $r$ values is not necessarily uniform. $n=160$ corresponds to the results without fixed spins. The annealing time is set to $\tau=2000.0~{}\mu s$. The ground state energy is $0.597$. Each panel represents a different error rate for fixing spins, with $p_{\mathrm{err}}$ set to (a) $p_{\mathrm{err}}=0.0$, (b) $p_{\mathrm{err}}=0.1$, (c) $p_{\mathrm{err}}=0.2$, (d) $p_{\mathrm{err}}=0.3$, (e) $p_{\mathrm{err}}=0.4$ and (f) $p_{\mathrm{err}}=0.5$, respectively.
@@
The normalization by $1/N$ ensures that $\mathcal{H}_{\mathrm{p}}$ is an extensive quantity. If the ground state is degenerate, determining the error rate becomes difficult because the ground state is not uniquely defined. Therefore, we consider the Ising model without degeneracy. The ferromagnetic Ising model with $h=0$ exhibits trivial twofold degeneracy. To eliminate this trivial degeneracy, we set $h=0.1$ and $J=1$. The ground state of this homogeneous fully connected Ising model is trivially the all-up state.
The D-Wave Advantage system 6.4 [38] was used as the quantum annealer for the experiment. This system features $5760$ physical qubits arranged in the Pegasus graph [39], ideally enabling the embedding of a complete graph with $177$ nodes through minor embedding [28, 29]. The chain length is determined by the graph topology of the qubits on hardware, depending on the number of couplers of each qubit. However, not all qubits in the system are functional. Consequently, the number of spins in the Ising model was set to $N=160$, the near maximum size that can be embedded in the quantum annealer.
The Ising model is embedded in the quantum annealer using minor embedding, and the chain strength—representing the interaction between physical qubits—must be adjusted to ensure that qubits within the same qubit chain align in the same direction. When we naively solve the problems in quantum annealer, the auto-scale method is employed [40]. In the algorithm of auto-scale method, the longer length of the qubit chain requires larger values of chain strength. However, the chain strength must remain within the range $[\mathcal{J}_{\mathrm{min}},\mathcal{J}_{\mathrm{max}}]$ for the interaction. For this study, the chain strength was set to the fixed value of $2.0$, the strongest value permissible in this configuration[40]. Therefore, the relative chain strength varies depending on the parameter range of the Ising model. When the parameter range of the Ising model is large, the chain strength becomes relatively small. Due to the stochastic behaviour, the ground state search in the quantum annealer was performed $100$ times, and the minimum energy from these runs was adopted as the quantum annealer’s energy output in this experiment.
Additionally, we investigated the energy gap using exact diagonalization for the same fully connected Ising model without considering the embedding effect. The energy gap analysis yielded similar results for the long annealing time limit. Finally, we investigated the system’s properties at the thermodynamic limit, accounting for the effects of rescaling and fixing spins.
The annealing schedule after rescaling is shown in appendix B.
@@ 3.1 Experiment on an actual quantum annealer
When embedding an Ising model with a large energy scale into a real quantum annealer, rescaling the energy scale becomes unavoidable. This rescaling is performed as described in Eq. (2). The experiment was conducted on the D-Wave Advantage system 6.4 to examine how the quantum annealer’s performance depends on the rescaling parameter.
Figure 2 illustrates the parameter dependence of the minimum energy among $100$ runs. Since the energy scale after rescaling differs from the pre-rescaling energy scale, the energy values are restored to the original Ising model, as expressed in Eq. (11), after solving them with the quantum annealer. The $x$-axis represents the number $n$ of spins after fixing spins, and the $y$-axis denotes the rescaling parameter $r$, as shown in Eq. (9). As $r$ decreases, the quantum annealer achieves lower energy, as shown in Fig 2 (a)-(c). This result indicates that a wide parameter range is necessary to maintain the performance of the quantum annealer.
@@
Equation (12) exhibits the maximum increase in the range of the local magnetic field due to fixed spins. Since we are dealing with a homogeneous fully connected ferromagnetic Ising model, after fixing spins with $p_{\mathrm{err}}=0$, the local magnetic fields correspond to this upper bound. When fixing spins with $p_{\mathrm{err}}=0$, the local magnetic fields change as follows:
$h^{\prime}=h+J(N-n).$ (13)
@@
The parameters associated with qubit chains are crucial for performing QA to the Ising model with dense interactions. Ideally, each qubit chain should correspond to a single logical spin, with all qubits within the chain aligned in the same direction. To ensure this, the chain strength should be set as large as possible. In this experiment, we fix the chain strength at its maximum value. However, even at maximum chain strength, qubits within a qubit chain may still fail to align uniformly, leading to broken qubit chains. High energy points observed in Fig. 2 (d)-(f) result from the breaking of the embedding chain in the hardware.
When the chain strength is fixed, the ratio of the local magnetic fields to the chain strength becomes a critical factor. The local magnetic field is distributed equally among the qubits within the qubit chain. Assigning sufficiently large values of the local magnetic fields is also an effective strategy for preventing the qubit chain from breaking. Therefore, the longer the chain length, the weaker local magnetic fields are assigned to each qubit. Consequently, as the spin length increases, the ratio of the local magnetic fields to the chain strength decreases. Therefore, with a small $h$, maintaining the qubit chain becomes increasingly difficult as its length grows.
When $n$ is large, the long qubit chain is required to express the Ising model in hardware. As a result, large $n$ leads to a small ratio of the local magnetic fields to the chain strength. In addition, increasing both $n$ and $p_{\mathrm{err}}$ results in a smaller $h^{\prime}$ from Eq. (14). Consequently, with large $n$ and $p_{\mathrm{err}}$, the qubit chain is more likely to break compared to other parameter ranges.
When $r$ is small, the ratio of the local magnetic fields to the chain strength becomes small, which renders the qubit chain susceptible to chain breaking. To set the larger ratio of the local magnetic fields to the chain strength is necessary to avoid chain break for small $r$. Therefore, expanding the parameter range is necessary to accommodate large-scale problems when embedding into the quantum annealer and preventing chain break. These results highlight the importance of fixing spins in the correct direction to achieve lower energy and preserve qubit chains on the hardware.
@@
Figure 4: (Color online) Color map of the minimum energy gap $\Delta_{\mathrm{min}}$ distribution in $n$ versus $r$ space for the homogeneous fully connected ferromagnetic Ising model with $N=160$. The spacing of $r$ values is not necessarily uniform. $n=160$ represents the results without fixing spins. Each panel corresponds to a different error rate $p_{\mathrm{err}}$ for fixing spins: (a) $p_{\mathrm{err}}=0.0$, (b) $p_{\mathrm{err}}=0.1$, (c) $p_{\mathrm{err}}=0.2$, (d) $p_{\mathrm{err}}=0.3$, (e) $p_{\mathrm{err}}=0.4$, (f) $p_{\mathrm{err}}=0.5$.
@@ 3.2 Energy gap analysis
Since the minimum energy gap between the ground state and the first excited state, $\Delta_{\mathrm{min}}$, reflects the performance of quantum annealing near the adiabatic limit condition [37, 41], we investigated its properties. Figure 3 shows $\Delta_{\mathrm{min}}$ of the Hamiltonian after fixing spins with $p_{\mathrm{err}}=0$ and rescaling with $r$. We perform exact diagonalization to obtain the minimum energy gap $\Delta_{\mathrm{min}}$.
The number of fixing spins $n$ at which $\Delta_{\mathrm{min}}$ is maximized varies for each $r$ as shown in Fig. 3. Energy gap $\Delta_{\mathrm{min}}$ increases with $n$, consistent with a previous study [21] without considering energy rescaling. However, the fixing spins method does not necessarily lead to an expansion of $\Delta_{\mathrm{min}}$ when rescaling is considered. For example, when $r=2.5$, the maximum $\Delta_{\mathrm{min}}$ appears at $n=140$. These results indicate a trade-off between the energy gap expansion due to fixing spins and the reduction of the energy gap caused by rescaling. Additionally, for $r=0.25$, $\Delta_{\mathrm{min}}$ decreases as $n$ increases. This demonstrates that the effects of fixing spins depend on the rescaling parameter. The findings emphasize the importance of exploring the relationship between fixing spins and rescaling.
Next, we investigated the behaviour of $\Delta_{\mathrm{min}}$, including errors in fixing spins. Figure 4 presents a heatmap of $\Delta_{\mathrm{min}}$ for different values of $r$ and $n$. Each panel corresponds to a different $p_{\mathrm{err}}$. As shown in Figure 4, as $p_{\mathrm{err}}$ increases or $r$ decreases, the optimal $n$ that maximizes $\Delta_{\mathrm{min}}$ becomes smaller. These results suggest that $r$ and $n$ must be carefully considered to achieve high-quality solutions. For $r=1.0$, fixing spins (lower $n$) improves $\Delta_{\mathrm{min}}$. In contrast, for $r=10$, fixing spins reduces $\Delta_{\mathrm{min}}$. These findings highlight the necessity of considering the rescaling parameters when fixing spins and adjusting parameters accordingly. Fixing spins has a positive effect at small $r$ but a negative effect at large $r$.
@@
$\Delta(s)$ =$=\left\{s^{2}\sin^{2}{\theta_{0}}\left(\frac{2h}{r}+\frac{1+\sin^{2}{\theta_{0}}}{s\sin{\theta}}\right)^{2}-\frac{(1+\sin^{2}{\theta_{0}})^{2}}{s^{2}\sin^{2}{\theta}}\right.$
$+\left.s^{2}\sin^{4}{\theta_{0}}\frac{4J^{2}}{r^{2}}+s[2\sin^{2}{\theta_{0}}-2(1-s)\cos^{3}{\theta_{0}}]\frac{2J}{r}\right.$
$\left.+4(1-s)^{2}\cos^{2}{\theta_{0}}\right\}^{\frac{1}{2}},\quad 0\leq s\leq 1,$ (15)
where $\theta_{0}$ is the rotation angle around the $y$-axis shown in Eq. (26). The angle $\theta_{0}$, which minimizes the energy, is numerically calculated.
@@
In Fig. 5, $\Delta_{\mathrm{min}}$ increases as $r$ decreases, in line with the increase in the local magnetic fields $h^{\prime}$ after fixing spins. This indicates that smaller values of $r$ are more favourable for quantum annealing, assuming no restrictions on the parameter range. The descending order of $h^{\prime}$ corresponds directly to the descending order of $\Delta_{\mathrm{min}}$. However, it is important to note that the allowed parameter range differs for each $r$ in Fig. 5, even though the parameter range is constrained in actual hardware. Therefore, the comparison is more advantageous for settings with smaller $r$.
We also investigated $\Delta_{\mathrm{min}}$ while considering the restrictions of the parameter range. The black lines in Fig. 5 illustrate a comparison of $\Delta_{\mathrm{min}}$ within the same $h$ range $[\mathcal{M}_{\mathrm{min}},\mathcal{M}_{\mathrm{max}}]$, where we assume $[-\mathcal{M},\mathcal{M}]$ with $|\mathcal{M}_{\mathrm{min}}|=|\mathcal{M}_{\mathrm{max}}|\equiv\mathcal{M}$. We focus solely on the $h$ range because, in this numerical simulation, the $h$ range is always broader than the $J$ range after fixing spins. According to Fig. 5, the point at which the maximum energy gap occurs in the line of $\mathcal{M}=0.8$ is at $r=1.0$. On the other hand, for the line of $\mathcal{M}=4.0$, the point where the maximum energy gap occurs is $r=2.0$. These results suggest that optimal rescaling parameters exist for different $h$ ranges. Therefore, the effects of fixing spins in the thermodynamic limit also depend on the rescaling parameter.
When the $h$ range is wide, the fixing spins method effectively expands $\Delta_{\mathrm{min}}$. In contrast, when the $h$ range is narrow, the fixing spins method becomes less effective because the rescaling effects, which reduce $\Delta_{\mathrm{min}}$, dominate over the benefits of expanding $\Delta_{\mathrm{min}}$ through fixing spins. These findings are also supported by the finite-size Ising model results in Fig. 4. The same considerations apply to the energy gap at finite sizes, even if the problem scale increases. Therefore, the properties of the energy gap with fixing spins and rescaling are extensive with respect to size.